OpenStudy (anonymous):

Simpligy by removing factors of 1. 1280v^10y^4 ________________ 100v^8y^2 simplified form is = ?

jimthompson5910 (jim_thompson5910):

First, reduce 1280 over 100 to get 128/10 = 64/5 Now subtract corresponding exponents: So for v it's 10-8= 2 and for y it's 4-2 = 2. These values will be the final exponents for the corresponding variables. This means that the final answer is \[\large \frac{64v^2y^2}{5}\]

OpenStudy (anonymous):

Thank you so much